OpenStudy (anonymous):

I'll give medals The distance traveled by a falling object is given by the formula d = 0.5gt^2 where d = distance, g = the acceleration due to gravity, and t = time. Solve this equation for g, and use your formula to determine the acceleration due to gravity if a baseball takes 10 seconds to hit the ground after being dropped from a height of 490 meters. Show all steps in your work.

OpenStudy (anonymous):

Just so you know this is a physics question so you should have posted it in physics

OpenStudy (anonymous):

@caters I was asked this question in my algebra class so I put it in the algebra section.

OpenStudy (anonymous):

The distance traveled by a falling object is given by the formula d = 0.5gt^2 where d = distance, g = the force of gravity, and t = time. Solve this equation for g, and use your formula to determine the force of gravity if a baseball takes 10 seconds to hit the ground after being dropped from a height of 490 feet. Show all steps in your work. NOTE: Actually, g is the gravitional acceleration - not a force. However, weight, which is caused by gravity is a force. d = 0.5gt^2 d/0.5t^2 = g g = d/0.5t^2 t = 10 sec d = 490 ft g = 9.8 ft/sec² ANSWER

OpenStudy (anonymous):

Thank you @callmevannah247

OpenStudy (anonymous):

Your Welcome!! @96Neko